4-(25) Find i(t) two ways: e) in the time domain 6) using the Laplace Transform + 10 cost Volts 12 lovolts 1H
Answer #1

ilt ti lo lost v ir 1o.v. m TH for tro kula 10+ 10 cost i (t) + 1. dilte dt o i (t) + di(t) dt 10+ Lo cost first order non hoj.et. 3(10+ lo cost) et de t. Schoe t + + lo et cost ) dt lo et + sloet 10 et wst dt Let t I - Si la e cost dt I - t bole. [+Io + 5 (sint + lost)+cet i et 10 oett 10 e (sint&lost) +C 2 ict.at t20 ilt=0 10 + 5 + c CE -15 ilt) 10 + 5 sin t+s cost-15e-t5 Lap lase method - from ean 10 + Lo wst eit) + dilt) dt taking Laplace both side I (8) + 8 IU) 10 + U ele set IB2 = lo + t 1118) n s2+1 tse (8+)) (0) (82 +1 ) Ico) a 208? + 10 (8 +)) (5) (x2+1) take Partial fraction- А 2082 to + A t coto 82+1 B sti15 si I(8) a ela Lo s + st 5 $2+1 s 82+1 take Laplace inverse both side # id) 114) 10 - 15 e eitt + 5 costt 5 sint ilt) = 10+